OpenStudy (anonymous):

three waiters at a local restaurant earned different amounts of money the past week. Jennifer earned $40 less than twice the amount Tony earned. Chris earned $60 more than Tony. How much did each waiter earn if they earned a total of $900?

OpenStudy (kkutie7):

jennifer=2(Tony)-40 Chris=Tony+60 Jennifer+Chris+Tony=900

OpenStudy (kkutie7):

900=(2(Tony)-40)+(Tony+60)+(Tony)

OpenStudy (kkutie7):

900=2Tony-40+Tony+60+Tony 900=2Tony+20+Tony+Tony 900=4Tony+20 880=4Tony 220=Tony

OpenStudy (kkutie7):

Jennifer=2(220)-40=400 Chris=220+60=280